OpenStudy (anonymous):

If x ^3 = -27, then what is x? tell me if i am wrong i think its 9

OpenStudy (anonymous):

would it be negative?

jimthompson5910 (jim_thompson5910):

replace x with 9 and see what happens x^3 = -27 9^3 = -27 729 = -27 .. this equation is false so x = 9 is not the answer

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

note: 9^3 = 9*9*9

jimthompson5910 (jim_thompson5910):

NOT 9^3 = 9*3

OpenStudy (anonymous):

no that 3 is supposed to be the tiny number in the corner

jimthompson5910 (jim_thompson5910):

yeah \(\LARGE x^3\) right?

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

that means x cubed

jimthompson5910 (jim_thompson5910):

when you cube a number, you have it multiplied 3 times by itself

jimthompson5910 (jim_thompson5910):

7^3 = 7*7*7 8^3 = 8*8*8 etc etc

OpenStudy (jdoe0001):

\(\large x ^{\color{red}{ 3}} = -27\implies \sqrt[{\color{red}{ 3}}]{x ^{\color{red}{ 3}}}=\sqrt[{\color{red}{ 3}}]{-27}\implies x=?\)

OpenStudy (anonymous):

i dont no anything about square roots

OpenStudy (anonymous):

oh i got the answer never mind @jdoe0001